OpenStudy (anonymous):

Please help :c The temperature of a chemical reaction ranges between 40°C and 180°C. The temperature is at its lowest point when t = 0, the temperature is 90 degrees and increasing. What is a sine function that would model this reaction?

OpenStudy (anonymous):

f(x) = 70 sin 12x + 110 f(x) = 110 sin 12x + 70 f(x) = 70 sin pi over 6x + 110 f(x) = 110 sin pi over 6x + 70

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Ugh this question is worded sooooo poorly :c I'm having trouble making sense of it...

OpenStudy (anonymous):

Me too.. I'm sorry this is the exact question verbatim :c

zepdrix (zepdrix):

Hmm let's see if we can figure it out :d

OpenStudy (anonymous):

mkay c: But how do we even start? ;-;

zepdrix (zepdrix):

At t=0, it's at 90 degrees and rising.